OpenStudy (anonymous):

A college basketball player makes 80% of his free throws. At the end of a game, his team is down one point. He is fouled attempting a three-pointer and hence awarded three free throws. Note that one free throw is worth one point. Assume each free throw is independent. What is the standard deviation of the number of free throws he would make?

OpenStudy (goformit100):

Take Probabily in use..

OpenStudy (anonymous):

I know missing all of them is 0.008 and i don't know how to get the probability for the rest

OpenStudy (goformit100):

You can do, I belive on you.

OpenStudy (anonymous):

hint? :(

jimthompson5910 (jim_thompson5910):

This is a binomial distribution problem where n = 3 (he attempts 3 free throws) p = 0.80 (the probability he makes any given free throw)

jimthompson5910 (jim_thompson5910):

the mean of the binomial distribution is xbar = n*p the standard deviation is sigma = sqrt(n*p*(1-p))

OpenStudy (anonymous):

How do i get the probability of making 1 free throw?

jimthompson5910 (jim_thompson5910):

by using the formula P(X = k) = (n C k)*(p)^(k)*(1-p)^(n-k) where in this case n = 3 p = 0.8 k = 1

OpenStudy (anonymous):

oh :/ didn't learn that yet

jimthompson5910 (jim_thompson5910):

you haven't learned about the binomial distribution yet?

OpenStudy (anonymous):

Nope, i don't know if we will